THE ROTATIONAL SPECTRA OF THE SILICON ISOTOPIC SPECIES OF SiCC

DAMIAN L. KOKKIN, \underline{CARL A. GOTTLIEB}, MICHAEL. C. McCARTHY, and PATRICK THAD-DEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138; and SANDRA BRÜNKEN, Laboratoire de Chimie Physique Moléculaire, École Polytechnique Fédérale de Lausanne, Station 6, CH-1015 Lausanne, Switzerland.

Until this work, the rotational spectra of the silicon isotopic species of SiCC were based almost entirely on astronomical frequencies, because only the fundamental 1_{0,1} − 0_{0,0} transition ^{29}SiC_{2} and ^{30}SiC_{2} had been measured in the laboratory.^{a} We have now derived precise rotational and centrifugal distortion constants from laboratory measurements of 35 transitions of each isotopic species between 140 and 360 GHz with J ≤ 10 and K_{a} ≤ 8. The rotational spectra calculated with the laboratory measured constants are about two orders of magnitude more accurate than that of He et al.,^{b} who determined the spectroscopic constants from about 20 lines of ^{29}SiC_{2} and of ^{30}SiC_{2} in the wide-line source IRC+10216. The new laboratory measurements should aid assignment of the silicon isotopic species of SiCC in the spectral line survey of IRC+10216 with the SMA,^{c} and in future observations with ALMA.

aR. D. Suenram, F. J. Lovas, and K. Matsumura, Astrophys. Journ. Lett. 342, L103 (1989)

bJ. H. He, Dinh-V-Trung, S. Kwok, H. S. P. Müller, Y. Zhang, T. Hasegawa, T. C. Peng, and Y. C. Huang, Astrophys. Journ.

Suppl. Ser., 177, 275 (2008).

cN. A. Patel, K. H. Young, S. Brünken, R. W. Wilson, P. Thaddeus, K. M. Menten, M. Reid, M. C. McCarthy, Dinh-V-Trung, C. A. Gottlieb, and A. Hedden, Astrophys. Journ., in press (2009).